OpenStudy (anonymous):

Could someone please help me better understand and solve Angular and Linear Velocity?

OpenStudy (experimentx):

linear velocity = angular velocity * radius of motion

OpenStudy (anonymous):

I've got something like 20yds radius, and a rate of revolution of 5 revolutions per 2 hours, I need to use angular velocity to change it to radians/day, and linear velocity to change it to ft/day. I have no clue where to start

OpenStudy (experimentx):

5 revolutions per 2 hours = 2\( \pi\) * 5/ (2* 60 * 60 s) = 2\( \pi\) * 5/ (2/24 day) ^^ angular velocity is above change yard into ft, and use top formula to get linear velocity.

OpenStudy (mertsj):